OpenStudy (anonymous):

The equation for the period of a pendulum is T=2 Square Root of L/g. How does your equation from 1B relate to the pendulum equation? Use the pendulum equation to calculate the period of a 1.50 m pendulum. Remember that the value of "g" is 9.81 m/s2. Compare your calculated period (from 3B) to your data by using the graph you constructed. Explain any possible sources for error if your graph does not agree with your calculated results.

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (irishboy123):

check your formula, then plug in numbers in your calculator.... \(\large T = \ 2 \pi \sqrt{\frac{L}{g}}\) and what are 1B and 3B?